OpenStudy (clamin):

An air force pilot must descend 1500 feet over a distance of 9000 feet to land smoothly on an aircraft carrier. What is the plane's angle of descent?

OpenStudy (crashonce):



OpenStudy (crashonce):

using the tan formula, tan angle = 1500/9000 solve from here please

OpenStudy (kittiwitti1):

Let's name this angle X.\[\angle X=\tan^{-1}{\frac{1500}{900}}\]

OpenStudy (crashonce):

@clamin

OpenStudy (kittiwitti1):

Oh, there was another solution up... sorry, didn't notice that.

OpenStudy (clamin):

i got .167

OpenStudy (crashonce):

in degrees?

OpenStudy (crashonce):

you should get about 59 degrees

OpenStudy (clamin):

you saaid 1500/9000

OpenStudy (kittiwitti1):

Wait, did you use my formula or Crash's solution -

OpenStudy (kittiwitti1):

Yeah, you have to continue on with the arctan.

OpenStudy (crashonce):

the methods are the same

OpenStudy (clamin):

idk how u got 59

OpenStudy (crashonce):

maybe u were in radians change it to degrees

OpenStudy (clamin):

i did

OpenStudy (kittiwitti1):

Don't you have to arctan the tan fraction to get the degrees?

OpenStudy (crashonce):

you have to use inverse tan, not tan

OpenStudy (kittiwitti1):

That's what I said ._.; Unless arctan doesn't mean inv tan. *googles*

OpenStudy (kittiwitti1):

Yup, arctan is inv tan.

OpenStudy (clamin):

i know how you got 59..you divide 1500/900 instead of 1500/9000

OpenStudy (kittiwitti1):

I got 9.46, and I double-checked to see that I'm in degree mode ._.;

OpenStudy (crashonce):

woops my bad

OpenStudy (kittiwitti1):

I just found an answer key and it says the answer's 9.5 degrees...

OpenStudy (kittiwitti1):

http://prntscr.com/6rvdpd

OpenStudy (anonymous):

1500/9000=0.1667 take atan(0.1667)=9.4 deg

OpenStudy (kittiwitti1):

9.5.\[\tan^{-1}\frac{1500}{9000}=9.462322208\approx 9.5\]

OpenStudy (anonymous):

.165 rad

OpenStudy (kittiwitti1):

^ yes.

OpenStudy (kittiwitti1):

Don't forget to close the question if you don't need anymore help :)